Pure acid is to be added to a 5% acid solution to obtain 95L of 60% solution. What amount of each should be used?
How many lilter of 100% pure acid should be used to make the solution?
How many liter of the 5% solution should be in the mixture?
Thank you

Answer by edjones(8007)

95-x * 1=95-x
x * .05=.05x
95 * .60=57
.
95-x+.05x=57
-.95x=-38
x=40 liters of the 5% solution should be in the mixture.
95-40=55 liters of 100% acid should be used to make the solution.
.
Ed